OpenStudy (btaylor):

Two identical 100 kg objects (of negligible size) are in space 100 m away from each other. The force of gravity causes both objects to accelerate towards one another. Find the time it takes for the objects to collide.

OpenStudy (cybershadow):

m1=m2 = 100 kg x = distance between the objects = 100m F = Gm1m2/x^2 therefore F = 6.6 x 10-11 * 100 * 100 / 100^2 = 6.6 x 10-11 also acceleration a = F/m = 6.6 x 10-11/100=6.6 x 10-9 so both objects are gonna move towards each other with an accelertaion a so they both will cover x/2 distance in same time and collide so we need to find that time in which each object covers a distance of x/2 s = ut + 1/2at^2 100/2 = 1/2 6.6 x 10-9 t^2 implies t^2 = 100/ 6.6 x 10-9 t = (150 * 10-8)^1/2 t = 12.2 * 10-4 seconds Keep me posted if it is right answer, I m not sure abt its correctness :)

OpenStudy (btaylor):

I don't believe that is right. Because they are constantly moving toward each other, acceleration is changing (always increasing).

OpenStudy (cybershadow):

eh yes , missed that point totally

OpenStudy (cybershadow):

ok wait it cud be solved by Integration or something F = Integrating x from 100 m to 0m (Gmm/x^2) as initial distance is 100 and when they collide x becomes 0 . Lemme Integrate

OpenStudy (btaylor):

The integral of the acceleration will yield velocity; the integral of the velocity will give the distance. I'm figuring it out, but I can't figure out how to account for both of them moving. We are only looking at the force on one. Should we treat the distance as 50?

OpenStudy (cybershadow):

yeah we shud treat the distance as 50 , as both the objects will move equal distance as they r movin with equal acceleration

OpenStudy (cybershadow):

got the soution yet?

OpenStudy (anonymous):

may be we can use conservation of energy, at any time where object have moved x \[-G\frac{m^2}{x_0}=2\frac{mv^2}{2}-G\frac{m^2}{x_0-2x}\] thus velocity of the object: \[v=\sqrt{Gm(\frac{1}{x_0-2x}-\frac{1}{x_0})}=\sqrt{GM\frac{2x}{x_0-2x}}\] then distance travelled in time 'dt' can be written as \[dx=vdt\] moving velocity to left side,integrating both sides \[\int_0^{x_0/2}\frac{dx}{\sqrt{GM\frac{2x}{x_0-2x}}}=\int_0^tdt\] does it make sense?

OpenStudy (anonymous):

in second line I missed one 'x_0' thus \[t=\frac{1}{\sqrt{GM}}\int_0^{x_0/2}\Big(\sqrt{\frac{x_0^2}{2x}-x_0}\Big)dx\] this is \[t=\frac{x_0}{\sqrt{GM}}\int_0^{x_0/2}\Big(\sqrt{\frac{1}{2x}-\frac{1}{x_0}}\Big)dx\]